### v1

@@ 5.1.2 Centralized Learning
The evaluation of model performance and sustainability across centralized configurations reveals critical insights into the trade-offs between predictive accuracy, energy efficiency, and overall sustainability. Figures 2 and 3 present the performance of various models using metrics such as NRMSE, MSE, RMSE, and MAE. The results show that model performance varies significantly depending on the selected metric. To ensure consistency and comparability, NRMSE was chosen as the primary metric, as it normalizes errors relative to the data scale, providing a more realistic and interpretable estimation.
@@
From the data in Tables 5.1.2 and 5.1.3, as well as the referenced figures, we analyzed the performance of the models for 7 and 10 timesteps. The findings reveal that the Leaky architecture is the only model that performs better with 7 timesteps. In contrast, all other models achieve superior performance and lower error rates with 10 timesteps, highlighting the importance of selecting an optimal timestep configuration for each model to balance computational efficiency and predictive accuracy.
@@
At first glance, the Sustainability Index used to evaluate the models does not reveal a straightforward trade-off. Reducing model complexity often leads to lower power consumption but can result in higher error rates, which negatively impacts the Sustainability Index (where a lower value of S indicates greater sustainability). Ultimately, the Sustainability Index reflects the balance between energy efficiency and performance, and the critical question becomes how much power consumption can be reduced without a significant sacrifice in predictive accuracy.
Among the centralized configurations, the Leaky Neuron stands out as the most sustainable model during the training process. Despite its higher prediction error compared to other models, its remarkably low power consumption compensates for this limitation, resulting in the lowest S. This makes the Leaky Neuron the best overall performer in terms of energy efficiency and environmental impact.
In contrast, the Alpha Neuron demonstrates the best predictive accuracy with 10 timesteps, indicating its ability to effectively handle the task. However, this performance comes at the expense of significantly higher power consumption, as the model operates as a computationally intensive filter. Consequently, it has the worst S among all evaluated models.
The Lapicque Neuron exhibits the poorest performance, with the highest RMSE and a power consumption higher than both the MLP and CNN models. This combination of low accuracy and high energy demand results in the lowest S, making the Lapicque Neuron the least favorable choice for centralized configurations.
The RLeaky model presents a particularly intriguing case. While it achieved its best performance at 10 timesteps with improved predictions and a lower NRMSE, this came at the cost of significantly higher power consumption during training. This trade-off highlights the challenge of balancing accuracy and energy efficiency, making the RLeaky model less suitable for applications where sustainability is a critical consideration.
The Echo State model demonstrates a better balance between accuracy and power consumption. Ranking second in terms of sustainability, it outperforms models such as the Lapicque Neuron, MLP, and CNN. Its architecture leverages energy efficiency without substantial performance compromises, making it a promising alternative for resource-constrained applications.
Finally, the Synaptic model, which aims to replicate the complex functions of brain cells, demonstrates significant drawbacks in terms of resource requirements. It consumes substantially more power and produces higher NRMSE compared to the MLP and CNN. This result suggests that the additional biological realism in the Synaptic model comes at the expense of computational efficiency and predictive accuracy, raising concerns about its practicality in real-world applications.
Collectively, these findings emphasize the importance of carefully balancing energy consumption, predictive performance, and sustainability when selecting models for specific tasks. While models like the Leaky Neuron offer an optimal trade-off for energy efficiency and environmental impact, others, such as the Alpha and RLeaky neurons, highlight the inherent trade-offs between accuracy and energy consumption. These insights provide valuable guidance for model selection in diverse application contexts.
@@
Centralized Cumulative results on NRMSE, Energy Consumption and Sustainability of considered models. Model NRMSE Consumption (Wh) $\mathcal{S}$ MLP 1.3465 0.44 25.401 CNN 1.3272 1.403 29.993 Echo State (RC) 1.42877 0.431 25.637 7 spikes/input 10 spikes/input 7 spikes/input 10 spikes/input 7 spikes/input 10 spikes/input Leaky 1.3168 1.3531 0.0742 0.0785 22.960 23.109 Lapicque 1.9429 1.7641 1.4044 1.797 32.414 33.377 RLeaky 1.3699 1.3055 1.4343 1.875 30.302 31.724 Synaptic 1.3896 1.42 1.3318 2.33 29.956 33.838 Alpha 1.2832 1.2751 2.0561 4.343 32.265 34.075
@@ 5.1.3 Federated Learning
The graphs in Figure 4 provide a comprehensive comparison of energy consumption, predictive performance, and sustainability of federated learning configurations across various machine learning models. In those graphs the number in the of the model name means the number of timesteps, e.g. the Leaky7 means the Leaky model with 7 timesteps.
The first graph illustrates the energy consumption of the models under the two paradigms, highlighting the efficiency advantage of centralized learning, which generally requires less energy due to localized computation. In contrast, federated learning models exhibit higher energy consumption, largely due to the communication overhead inherent in decentralized training.
The second graph compares the Normalized Root Mean Square Error (NRMSE) of the models, offering insight into their predictive accuracy. Centralized models typically achieve better performance, reflected in their lower NRMSE values, as they can directly access and process all data. Federated models, while exhibiting higher error rates, still maintain reasonable predictive performance, making them suitable for scenarios where privacy and decentralization are prioritized.
The third graph examines the trade-off between energy consumption and NRMSE, visualizing the balance between resource efficiency and predictive accuracy. This trade-off is critical for assessing the suitability of different learning paradigms in energy-constrained environments, such as edge devices or mobile applications. Federated learning’s higher energy costs and slightly reduced accuracy underscore the need for optimization strategies to improve its feasibility in such contexts.
The final graph compares the S across models, integrating both energy efficiency and predictive performance. Centralized models generally achieve higher sustainability due to their lower energy requirements and superior accuracy. Federated models, however, show varying levels of sustainability depending on the architecture, emphasizing the importance of selecting the appropriate paradigm based on application-specific trade-offs among energy, accuracy, and sustainability.
Additionally, from Table 5.1.3, we observe that in the federated setup, the CNN model achieves the most accurate predictions. Its performance is comparable to that of the neural network architecture with Alpha neurons, which outperform the traditional MLP model.
Furthermore, predictions made with Lapicque neurons are significantly less accurate, with an NRMSE approximately twice as high as that of the CNN model.
The Leaky neuron also performs competitively in the federated setup, achieving a lower NRMSE than the MLP, although its error remains slightly higher than that of the CNN.
Figure 4: Federated Learning Results
@@
Federated Cumulative results on NRMSE, Energy Consumption and Sustainability of considered models. Model NRMSE Consumption (Wh) $\mathcal{S}$ MLP 3.3134 0.28 29.86 CNN 2.9999 0.64 32.410 Echo State (RC) 4.9471 0.105 36.034 7 spikes/input 10 spikes/input 7 spikes/input 10 spikes/input 7 spikes/input 10 spikes/input Leaky 3.3061 3.4012 0.0138 0.0213 27.6385 27.9062 Lapicque 6.141 5.131 0.751 0.939 39,1163 38,4670 RLeaky 3.332 3.101 0.832 0.812 33,6647 32.9419 Synaptic 3.621 3.231 0.891 1.234 34,7515 35,6637 Alpha 3.2813 3.1023 1.902 2.523 39.0313 41,0282
@@
From Figures 5 and 6, each contain 4 graphs, we observe that in our simulation, federated power consumption is significantly lower compared to centralized configurations. However, this reduction in power consumption comes at the cost of a noticeably higher NRMSE. This trade-off highlights a fundamental challenge in federated learning, where decentralized training introduces communication overheads and reduced data accessibility, leading to diminished predictive accuracy.
Furthermore, for the federated configuration, we find that all models, except the Leaky model, achieve better NRMSE performance with 10 timesteps. In contrast, the Leaky model demonstrates improved metrics with 7 timesteps. This observation underscores the importance of selecting an optimal timestep configuration to balance computational complexity, energy efficiency, and predictive accuracy.
The first graph illustrates the energy consumption of centralized and federated models. Centralized configurations demonstrate consistently lower energy consumption due to localized computations, whereas federated models exhibit higher energy usage owing to the communication overhead inherent in decentralized training.
The second graph compares the predictive performance of the models using the Normalized Root Mean Square Error (NRMSE) metric. Centralized models achieve superior accuracy, reflected in their lower NRMSE values, due to their direct access to the entire dataset. Federated models, while showing higher NRMSE values, maintain reasonable performance levels, making them viable for applications where privacy and decentralization are prioritized over predictive precision.
The third graph visualizes the trade-off between energy consumption and predictive accuracy, with a scatter plot of energy consumption versus NRMSE. Points closer to the origin (0, 0) indicate better resource efficiency and lower predictive error. Centralized models cluster near the origin, showcasing superior energy efficiency and predictive accuracy, while federated models demonstrate a clear trade-off with higher energy costs and slightly reduced accuracy.
The final graph compares the sustainability index (S), which integrates energy efficiency and predictive performance. Centralized models achieve higher sustainability scores due to their lower energy consumption and better accuracy. Federated models exhibit varying levels of sustainability, depending on the architecture and configuration, underscoring the importance of evaluating application-specific trade-offs between energy, accuracy, and sustainability.
@@
We observe that energy consumption for federated learning is lower in our simulations than centralized, but the NRMSE values are higher. This indicates that centralized models are more sustainable, as they achieve a lower Sustainability Index (S) compared to federated models. This observation is further supported by the Sustainability Index Comparison Graph, which directly compares the S values of federated and centralized configurations.
In the Energy vs. NRMSE Trade-off graph from the same figures, the blue dots (representing centralized models) exhibit much lower NRMSE values. In this graph, dots closer to the origin (0, 0) represent models with lower S, given the constant dataset size. It is evident from this analysis that the most sustainable models across all experiments are the centralized Leaky models with 7 timesteps and 10 timesteps. Among these, the centralized Leaky model with 7 timesteps is the most sustainable, achieving an S value of 22.960.
Figure 5: Federated vs Centralized Learning 7 timesteps.
Figure 6: Federated vs Centralized Learning 10 timesteps.
@@ 6 Discussion and Open Issues
We observe significant variations in power consumption across models during training. For example, the Alpha Neuron architecture for SNNs achieves lower prediction errors but at the cost of substantially higher power consumption. Notably, it demonstrates the best predictive capacity in our centralized training experiments.
@@
These findings highlight that there is no universal rule ensuring bio-inspired models inherently consume less power while maintaining low predictive error. Consequently, a clear trade-off cannot be universally established; each case must instead be analyzed individually.
@@
We do, however, observe a general trend: increased computational complexity tends to result in higher power consumption alongside improved predictive performance. The challenge, therefore, is to select a model with a level of computational complexity that balances acceptable prediction error and manageable power consumption for a specific application.
In our application—cellular traffic forecasting—this consideration is especially critical. The vast amounts of data generated daily raise concerns about significant CO2 emissions resulting from large-scale computations.
@@
At the same time, we must acknowledge the limitations of current models in accurately simulating brain-like learning mechanisms. While the Alpha neuron achieves better predictive capacity by incorporating spiking elements that mimic brain functionality, its internal structure resembles a filter more than the biological structure of a neuron. This raises an intriguing question: how does the human brain achieve such remarkable learning capabilities with such low power consumption?
This question remains unresolved. Meanwhile, the Synaptic neuron, which more closely resembles the functionality of brain cells, exhibits higher error rates compared to classic models but performs better than the Alpha neuron. Additionally, the Synaptic neuron consumes significantly more power than classic models, though less than the Alpha neuron.
@@
The increased power consumption of the Synaptic neuron arises from its attempt to simulate biological processes more accurately, resulting in more complex computations. This suggests there may be a limit to how closely we should model brain functionality. In our case, the more biologically accurate Synaptic neuron produced more erroneous results compared to the Leaky neuron, a simplified, classic Leaky-Integrate-and-Fire neuron.
These observations lead to two important considerations:
  1. 1.
    Specialized versus general-purpose hardware: The cells in our brain are specialized to perform specific functions, whereas our experiments are conducted on general-purpose hardware. This underscores the need to replicate these experiments on neuromorphic hardware designed specifically for each model. Such hardware would allow us to fully exploit their advantages while minimizing disadvantages. Furthermore, specialized hardware would provide more accurate estimations of power consumption, as computational simulations are prone to distortion.
  2. 2.
    Task alignment with brain-like models: Tasks involving numerical pattern recognition and approximate prediction may not align well with the human brain’s natural capabilities. It may be unreasonable to expect a model designed to mimic the human brain to excel in such tasks without question. This does not diminish the learning power of the human brain but instead suggests that we should focus on identifying tasks where these models perform best, reflecting the strengths of human brain functionality.
In addition, the power consumption reported for federated learning in our experiments does not account for the energy required for communication, as these experiments were conducted in a simulated environment. This limitation highlights the need for further studies incorporating real-world network conditions. Such studies should evaluate the true energy footprint of federated learning, particularly in scenarios with varying communication overheads and device heterogeneity.

### v2

@@ 5.1.2 Centralized Learning
The evaluation of model performance and sustainability across centralized configurations reveals critical insights into the trade-offs between predictive accuracy, energy efficiency, and overall sustainability. Figure 1 present the performance of various models using metrics such as NRMSE, MSE, RMSE, and MAE. The results show that model performance varies significantly depending on the selected metric. To ensure consistency and comparability, NRMSE was chosen as the primary metric, as it normalizes errors relative to the data scale, providing a more realistic and interpretable estimation.
@@
From the data in Tables 5.1.2 and 5.1.3, as well as the referenced figures, we analyzed the performance of the models for 7 and 10 timesteps. The findings reveal that the Leaky and RLeaky architectures are the models that perform better with 7 timesteps. In contrast, all other models achieve superior performance and lower error rates with 10 timesteps, highlighting the importance of selecting an optimal timestep configuration for each model to balance computational efficiency and predictive accuracy.
@@
Regarding sustainability, an initial inspection of the Sustainability Index suggests that the relationship between energy efficiency and predictive performance is not straightforward. Reducing model complexity often leads to lower power consumption but can result in higher error rates, which negatively impacts the Sustainability Index (where a lower value of S indicates greater sustainability). Ultimately, the Sustainability Index reflects the balance between energy efficiency and performance, and the critical question becomes how much power consumption can be reduced without a significant sacrifice in predictive accuracy. Among the centralized configurations, the Leaky Neuron stands out as the most sustainable model during the training process. Despite its higher prediction error compared to RNN, its predictive capacity compared to MLP, CNN, and its remarkably low power consumption compensates for this limitation, resulting in the lowest S. This makes the Leaky Neuron the best overall performer in terms of the examined trade-off.
In contrast, the Alpha Neuron with 7 timesteps demonstrates the best predictive accuracy among all centralized models, indicating its ability to effectively handle the task. However, this performance comes at the expense of significantly higher power consumption, as the model operates as a computationally intensive filter. Consequently, it has the worst S among all evaluated models. The Lapicque Neuron exhibits the poorest performance, with the highest RMSE and a power consumption higher than both the MLP and CNN models. This combination of low accuracy and high energy demand results in the lowest S, making the Lapicque Neuron the least favorable choice for centralized configurations. The RLeaky model presents a particularly intriguing case. While it achieved its best performance at 10 timesteps with improved predictions and a lower NRMSE, this came at the cost of significantly higher power consumption during training. This trade-off highlights the challenge of balancing accuracy and energy efficiency, making the RLeaky model less suitable for applications where sustainability is a critical consideration. The Echo State model demonstrates a better balance between accuracy and power consumption. Ranking second in terms of sustainability, it outperforms models such as the Lapicque Neuron, MLP, and CNN. Its architecture leverages energy efficiency without substantial performance compromises, making it a promising alternative for resource-constrained applications. Finally, the Synaptic model, which aims to replicate the complex functions of brain cells, demonstrates significant drawbacks in terms of resource requirements. It consumes substantially more power and produces higher NRMSE compared to the MLP and CNN. This result suggests that the additional biological realism in the Synaptic model comes at the expense of computational efficiency and predictive accuracy, raising concerns about its practicality in real-world applications.
Collectively, these findings highlight that certain bio-inspired models are capable of reducing power consumption, which in turn contributes to lower carbon emissions.
@@
Centralized Cumulative results on NRMSE, Energy Consumption and Sustainability of considered models. Model NRMSE Consumption (Wh) $\mathcal{S}$ MLP 1.3465 0.44 25.398 RNN 1.2693 1.624 30.62 CNN 1.3272 1.403 29.992 Echo State (RC) 1.3343 0.431 25.302 7 spikes/input 10 spikes/input 7 spikes/input 10 spikes/input 7 spikes/input 10 spikes/input Leaky 1.2870 1.2864 0.0531 0.0713 22.713 22.839 Lapicque 1.6099 1.6358 1.5143 1.813 31.6174 32.9177 RLeaky 1.2732 1.2678 1.5343 1.862 30.287 31.503 Synaptic 1.2832 1.3132 1.3623 2.543 29.6362 34.024 Alpha 1.2639 1.2872 2.2141 3.234 32.713 35.95
@@ 5.1.3 Federated Learning
The graphs in Figure 2 provide a comprehensive comparison of energy consumption, predictive performance, and sustainability of federated learning configurations across various machine learning models. In those graphs the number in the model name means the number of timesteps, e.g. the Leaky7 means the Leaky model with 7 timesteps.
The first graph illustrates the energy consumption of the models under the two paradigms, 7 and 10 timesteps, giving us a point to understand how in this federated setup the power consumption, reflects the predictive capacity. The second graph compares the Normalized Root Mean Square Error (NRMSE) of the models, offering insight into their predictive accuracy. The third graph examines the trade-off between energy consumption and NRMSE, visualizing the balance between resource efficiency and predictive accuracy. This trade-off is critical for assessing the suitability of different learning paradigms in energy-constrained environments, such as edge devices or mobile applications. The final graph compares the S across models, integrating both energy efficiency and predictive performance. Federated models, however, show varying levels of sustainability depending on the architecture, emphasizing the importance of selecting the appropriate paradigm based on application-specific trade-offs among energy, accuracy, and sustainability.
Additionally, from Table 5.1.3, we observe that in the federated setup, the RNN model achieves the most accurate predictions. Its performance is higher comparable to that to the next better neural network architecture with Alpha neurons. Alpha outperform the traditional MLP and CNN model. Furthermore, predictions made with Lapicque neurons are significantly less accurate, with an NRMSE approximately twice as high as that of the CNN model. The Leaky neuron also performs competitively in the federated setup, achieving a lower NRMSE than the MLP, although its error remains slightly higher than that of the CNN and RNN.
Figure 2: Federated Learning Results (i) the Power consumption (Green) and the NRMSE (Pink) across different models in federated learning. (ii) the sustainability index (S) for the federated models. (iii) The trade-off between the NRMSE and the Power Consumption, the color of each dot represents the sustainability index. (iv) The sustainability index across different models, starting the y axis from 0.
@@
Federated Cumulative results on NRMSE, Energy Consumption and Sustainability of considered models. Model NRMSE Consumption (Wh) $\mathcal{S}$ MLP 3.3134 0.28 29.86 RNN 1.529 0.832 28.186 CNN 2.9999 0.64 32.410 Echo State (RC) 4.8372 0.096 31.3536 7 spikes/input 10 spikes/input 7 spikes/input 10 spikes/input 7 spikes/input 10 spikes/input Leaky 3.3019 3.3854 0.0125 0.0207 27.6385 27.8677 Lapicque 5.869 5.127 0.728 0.895 38.4474 38.1685 RLeaky 3.289 3.046 0.842 0.890 33.6144 33.2547 Synaptic 3.598 3.126 0.945 1.165 35.0182 35.0049 Alpha 3.2765 3.0984 1.765 2.265 38.3991 39.9988
@@
From Figure 3, which comprises four visualizations, we observe that in our simulation, federated power consumption is significantly lower compared to centralized configurations. However, this reduction in power consumption comes at the cost of a noticeably higher NRMSE. Furthermore, for the federated configuration, we find that for different models and configurations the performance preserves a volatility with 7 and 10 timesteps. This observation underscores the importance of selecting an optimal timestep configuration to balance computational complexity, energy efficiency, and predictive accuracy.
The first graph of figure 3 illustrates the energy consumption of centralized and federated models. Federated configurations demonstrate consistently lower energy consumption, whereas centralized models exhibit higher energy usage. The second graph compares the predictive performance of the models using the Normalized Root Mean Square Error (NRMSE) metric. Centralized models achieve superior predictions, reflected in their lower NRMSE values. Federated models, while showing higher NRMSE values, maintain reasonable performance levels, making them viable for applications where energy restrictions exists and power consumption are prioritized over predictive precision. The third graph visualizes the trade-off between energy consumption and predictive accuracy, with a scatter plot of energy consumption versus NRMSE. Points closer to the origin (0, 0) represent models that achieve both high resource efficiency and low predictive error. Centralized models tend to cluster near the lower end of the NRMSE axis, indicating higher accuracy, but many of them exhibit higher values along the power consumption axis. In contrast, federated models demonstrate a clearer trade-off: most show lower energy consumption but at the cost of slightly reduced predictive accuracy.
The final graph compares the sustainability index (S), which integrates energy efficiency and predictive performance. Centralized models achieve better sustainability scores due to their higher predictive capacity. Federated models exhibit varying levels of sustainability depending on the architecture and configuration; however, they generally underperform compared to centralized configurations in terms of sustainability.
@@
It is evident from this analysis that the most sustainable models across all experiments are the centralized Leaky models with 7 timesteps and 10 timesteps. Among these, the centralized Leaky model with 7 timesteps is the most sustainable, achieving an S value of 22.713. In other words, this model achieves the best overall trade-off among energy consumption, predictive accuracy, and communication cost. However, this outcome can vary depending on the specific priorities of a given application. By adjusting the exponents in the Sustainability Index, one can emphasize particular aspects of the trade-off to better align with application-specific requirements.
Figure 3: Federated vs Centralized Learning 7 timesteps. (i) Comparison of the Power Consumption of the centralized (Purple) and the federated (yellow) (ii) Comparison of the NRMSE of the centralized (Purple) and the federated (yellow) (iii) Energy and NRMSE trade off, the color indicates the federated or the centralized configuration, and the radius of the circle the magnitude of the sustainabilitay index. (iv) Comparison of the Sustainability index (S) of the centralized (Purple) and the federated (yellow).
@@ 6 Discussion and Open Issues
We observe significant variations in power consumption across models during training. For example, the Alpha Neuron architecture for SNNs achieves lower prediction errors but at the cost of sustainability, due to higher power consumption. Notably, it demonstrates the best predictive capacity in our centralized training experiments. These findings highlight that there is no universal rule ensuring bio-inspired models inherently consume less power while maintaining low predictive error. Consequently, a clear trade-off cannot be universally established; each case must instead be analyzed individually.
@@
We do, however, observe a general trend: increased computational complexity tends to result in higher power consumption alongside improved predictive performance. The challenge, therefore, is to select a model with a level of complexity that balances acceptable prediction error with manageable energy demands, tailored to the needs of a specific application.
@@
At the same time, we must acknowledge the limitations of current models in accurately simulating brain-like learning mechanisms. While the Alpha neuron achieves better predictive capacity by incorporating spiking elements that mimic brain functionality, its internal structure resembles a filter more than the biological structure of a neuron.
@@
Meanwhile, the Synaptic neuron, which more closely resembles the functionality of brain cells, exhibits higher error rates compared to classic models but performs better than the Alpha neuron. Additionally, the Synaptic neuron consumes significantly more power than classic models, though less than the Alpha neuron. The increased power consumption of the Synaptic neuron arises from its attempt to simulate biological processes more accurately, resulting in more complex computations. This suggests there may be a limit to how closely we should model brain functionality. In our case, the more biologically accurate Synaptic neuron produced more erroneous results compared to the Leaky neuron, a simplified, classic Leaky-Integrate-and-Fire neuron.
These observations raise two key points: First, unlike the brain’s specialized cells, our models run on general-purpose hardware, highlighting the need to evaluate them on neuromorphic systems that better reflect their design and provide more accurate power measurements. Second, brain-inspired models may not naturally align with tasks like numerical prediction, so their strengths should be assessed in domains that better match biological processing capabilities.
In our experiments, centralized learning most outperformed federated learning in terms of predictive accuracy and the energy-prediction trade-off. For example, the centralized Echo State Network (ESN) achieved an NRMSE of 1.3343 and consumed 0.431 Wh, resulting in a sustainability index S = 25.302. In contrast, the federated ESN yielded a much higher NRMSE of 4.8372 despite consuming less energy (0.096 Wh), with a significantly worse sustainability index S = 31.3536. Similarly, the centralized Leaky model (7 timesteps) achieved the best overall sustainability (S = 22.713) with NRMSE = 1.2870, outperforming its federated counterpart (S = 27.6385, NRMSE = 3.3019) despite slightly higher energy usage (0.0531 Wh vs. 0.0125 Wh). The RNN model also demonstrates strong predictive performance on our time-series data in both centralized and federated settings, confirming its effectiveness for temporal tasks. However, its high energy consumption limits its suitability to resource-rich environments. Notably, the federated RNN achieved an NRMSE = 1.529, outperforming all other models in the federated setting.
Finally, it is worth highlighting some key aspects regarding the sustainability analysis.
In the context of this work, federated learning proved more energy-efficient during the model training phase. A possible explanation of this is the early stopping mechanism applied at each client, which led to faster model convergence and reduced local computations in the decentralized setting.
Another important sustainability factor is data transmission. As previously discussed, federated learning requires only minimal data exchange—typically a few kilobytes of model weights—thereby improving the overall sustainability index.
Despite these advantages, federated models consistently underperformed in terms of prediction error across all experiments. This performance gap contributes to centralized training achieving a better overall Sustainability Index, as the metric also accounts for model accuracy.
Potential future directions include exploring hybrid architectures that combine SNNs with conventional RNNs and ESNs, as well as a variety of ensembling strategies—from simple vote-based schemes to component-level integrations—to assess their trade-offs between predictive performance and power efficiency within a sustainability framework. However, these approaches require careful consideration, as simply executing multiple models and aggregating their outputs can substantially increase power consumption and undermine sustainability goals. A very interesting future work will be to experiment with different encoding techniques of the input to timesteps. Also, we aim to experiment with alternative architectures for Reservoir Computing, as well as explore various nature-inspired optimization algorithms with a potential to reduce the power consumption —such Intelligent Water Drops as it described in the Siddique et al.  [30], Bacterial Foraging as it described in Wand et al.  [31], and Spiral dynamics algorithms as described in  [32] potential enhancements in the next phase of our research.